OpenStudy (anonymous):

PLEASE HELP I WILL FAN AND MEDAL! The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1150. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many one-step equations eliminated today? 430 720 1150 1300

OpenStudy (anonymous):

Is the correct answer 720?

OpenStudy (anonymous):

@Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

one-step equation will be represented by \(O\) and two-step equations will be represented by \(T\)

HanAkoSolo (jamierox4ev3r):

if you follow what the problem says, 3\(O\)-2\(T\)=1300

OpenStudy (anonymous):

the answer is 720.

HanAkoSolo (jamierox4ev3r):

@FilthyMcNasty i was in the middle of explaining... -.-

HanAkoSolo (jamierox4ev3r):

solve for each variable and you will indeed get 720. This means that you are correct @aleximhappy08

OpenStudy (anonymous):

x+y=1150 3x-2y=1300 y=-x+1150 3x-2(-x+1150)=1300 3x+2x-2300=1300 5x-2300=1300 +2300 +2300 5x=3600 x=720

OpenStudy (anonymous):

Thanks!

HanAkoSolo (jamierox4ev3r):

i chose the variables O and T, but you could also use x and y if you want to make things easier. and no problem :)

OpenStudy (anonymous):

:)